# Biased Minds Meet Biased AI: How Class Imbalance Shapes Appropriate Reliance and Interacts with Human Base Rate Neglect

### v1

@@ Abstract.
Figure 1. Conceptual illustration of AI class imbalance, its implications for appropriate reliance calibration and compounding effects with human base rate neglect.:
@@ 1. Introduction
Artificial intelligence (AI) is increasingly present in daily life, raising concerns about biased outputs and their societal implications (O’neil2016Weapons; Mallari2020criminal). In the human-computer interaction (HCI) community, there is a large body of research trying to make AI less biased and therefore more reliable, fair, and transparent (Adadi2018Peeking; Lepri2018Fair). At the same time, we have a long and deep understanding that human decision-making is susceptible to cognitive biases, causing decisions to deviate from rationality (Tversky1974Judgment; Wilke2012Cognitive). As human-AI interaction becomes more common, it is critical to understand what happens when biased AI decision-support meets biased human decision-making. This interplay between human and AI biases has been termed compound human-AI bias (vonfeltenIsolationInteractionistPerspective2025).
@@
One possible outcome of such compounding is bias amplification, as observed in areas like emotion perception, motion discrimination, and social judgments (glickmanHowHumanAI2024), or investment decisions (winderBiasedEchoesLarge2025). However, while bias amplification has been studied, the broader dynamics of bias interactions are still not well understood. In particular, the absence of AI bias may also mitigate human bias, suggesting that AI systems have the potential not only to exacerbate existing prejudices but also to influence human decision-making in a positive direction. Understanding these interactions requires examining how human and AI biases influence one another, as well as the conditions under which AI may serve as a corrective mechanism rather than a reinforcing agent of bias. In human-AI interaction, decisions often manifest in reliance, making it a critical outcome for examining how human and AI biases jointly shape behaviour.
@@
Reliance can be defined as a “user’s behavior that follows from the advice of the system” (Scharowski2022Trust, p. 3). It must be calibrated to AI’s performance to ensure appropriate reliance. Appropriate reliance means relying on AI when it is correct and disregarding it when it is incorrect  (Lee2004Trust; Schemmer2023Appropriate). However, AI bias could complicate and distort this calibration. A common source of bias in AI is class imbalance, which occurs when certain outcome variables are under-represented in AI training data (Fazelpour2021Algorithmic). This imbalance leads to AI systems favouring the majority class, reducing prediction accuracy on new data. While AI bias can skew the AI’s predictions, human biases could further complicate AI-assisted decision-making.
@@
A common human bias, base rate neglect (BRN), can be defined as the tendency to underweight base rate information when estimating probabilities (Yang2020Base). It poses challenges in decision-making, particularly in AI-assisted tools (babic2021direct). Although many AI tools are developed specifically for experts (Panigutti2022understanding), non-expert users frequently rely on applications such as symptom checkers or disease screening tools (Cheng2019Explaining), where BRN is important (babic2021direct). Erroneous decisions caused by biases, such as BRN, could increase user distress (ryen2009cyberchondria) and translate to systemic outcomes like increased healthcare costs (Chambers2019Digital). Despite research on technical AI bias mitigation (Deshpande2020Mitigating; Ramaswamy2021Fair; Ranaldi2023Trip; Zhou2022Bias), as well as a plethora of research on cognitive bias mitigation (kortelingRetentionTransferCognitive2021), little attention has been given to the biases of humans when interacting with AI. Human-centred AI research is only recently highlighting the importance of cognitive biases in human-AI interactions (Boonprakong2025cogbiases; Kliegr2021review; Bertrand2022How; Wang2019Designing; liao2021xai). However, experimental studies on these interactions are scarce, and the compound effects of AI and human biases remain underexplored.
@@ 2.1. AI Biases: Class Imbalance
Algorithmic bias, or AI bias, refers to systematic deviations in AI outputs caused by biased data at different stages of development, such as data collection, modelling, or deployment (Fazelpour2021Algorithmic). Since AI learns from patterns in its training data, biased data can lead to biased outcomes. These biases may stem from societal inequalities (e.g., gender bias in medical diagnoses (AlHamid2024)) or from decisions during data collection that create unrepresentative datasets (Fazelpour2021Algorithmic). For instance, training a model on data that underdiagnoses heart disease in women could lead to male-centric predictions, ultimately reinforcing healthcare disparities (AlHamid2024). A common form of data bias is sampling bias, where certain groups are underrepresented, leading to predictions that reinforce existing imbalances (Srinivasan2021Biases). Consider a disease screening tool trained on data collected during a dengue outbreak, where dengue cases were more prevalent than influenza or the common cold. Due to this class imbalance, the tool may later overdiagnose dengue while neglecting the more common diseases.
@@ 2.2. Human-AI Decision-Making: Reliance
Figure 2. Illustration of reliance outcome Combinatorics. This figure was adapted from Schemmer2023Appropriate.:
Human-AI decision-making, also referred to as AI-assisted decision-making, refers to a collaborative process between humans and AI where AI advice supports users in forming final judgments or decisions (Chen_2023). A common objective in this context is determining whether the AI advice helps users make better decisions (e.g., nourani_anchoring_2021; Liu_2021; Zhang2022complementary). For instance, an AI-based screening tool for diagnosing dengue fever cases should ideally improve diagnostic accuracy beyond the user’s judgment or the AI’s advice alone.
In this collaborative process, reliance is a critical concept for assessing the effectiveness of human-AI interaction. Reliance can be quantified by assessing how often or how much users adjust their behaviour after receiving AI advice. However, reliance is only beneficial if it is calibrated to the AI’s capabilities (Lee2004Trust). Ideally, users should demonstrate appropriate reliance (Schemmer2023Appropriate; He2023Dunning), relying on the AI when it is correct and refraining from relying when it is incorrect. To measure reliance in these scenarios, a step-wise evaluation of AI advice in the presence of systematic errors is necessary. Collaborative decisions must be compared against uninfluenced human decisions as a baseline. Following these considerations, Schemmer2023Appropriate proposed a measurement concept for classification problems to assess the appropriateness of reliance, distinguishing four reliance outcomes:
@@
Appropriateness of Reliance is captured by two measures: relative self-reliance (RSR) and relative AI-reliance (RAIR). Optimally, both would equal 1, meaning users consistently rely on themselves when they are correct, and the AI is not (RSR=1), and on the AI when it is correct, and they are not (RAIR=1). Together, these measures form a two-dimensional framework, which enables us to test how AI bias affects dimensions of appropriate reliance. However, beyond appropriate reliance, AI bias may also temporally interact with human biases when shaping decisions. If users interact with biased AI, their own biases may compound errors made by the AI, creating a feedback loop (glickmanHowHumanAI2024).
@@ 2.3. Human Biases: Base Rate Neglect
To understand the phenomenon of BRN, a fundamental understanding of Bayes’ theorem is essential. Bayes’ theorem helps to understand how probabilities change after observing new evidence. For example, it helps us determine the probability that a patient has a disease after a positive test result. More formally, it calculates the conditional probability of a disease $H_{\text{disease}}$ given the data $D_{\text{test positive}}$ (for a comprehensive explanation see (webb_bayes_2020)). A well-known example illustrating the importance of base rates is regular breast cancer screening for women over 40 (Eddy1982Probabilistic). Even with mammograms showing 80% sensitivity and 90% specificity, the low base rate of breast cancer (1%) means that only about 7% of women who test positive once actually have cancer. This probability, known as the positive predictive value (PPV), often surprises people, including physicians, who mistakenly report sensitivity or the difference between sensitivity and specificity as PPV (Gigerenzer1998AIDS; Hoffrage1998Using). Because of this neglect of base rates, the usefulness of such screenings is still debated in contemporary medical literature (Jørgensen2016Breast). Traditionally, experiments have focused on assessing BRN through deviations from PPV, requiring participants to integrate base rates with sensitivity and specificity (Dunwoody2005Use). However, this approach may even evoke BRN by overwhelming participants with complex statistical concepts (Todd2002Testing). Patients were shown to mistake sensitivity for PPV in common diseases like strep throat and heart attacks (hamm_accuracy_1998), and even expert clinicians face challenges with these judgments (Gigerenzer1998AIDS; gigerenzer_helping_2007). To address challenges in probability judgments, the use of natural frequencies (e.g., one out of 1000) has been proposed as an alternative to relative frequencies (e.g., 0.1%) to reduce complexity and mitigate BRN (Hoffrage2002Representation). However, even natural frequency formats still require some mental calculations. We argue that to assess BRN with less complexity for participants, presenting them verbally (e.g., ”very often,” ”rarely”) could be more effective. This eliminates the need for mental calculations by participants but still allows to infer BRN by comparing decisions between low- and high-base rate disease. A further advantage is that it does not require participants to know the true population base rate.
@@ 2.4. Human-AI Biases in AI-Assisted Decision-Making
AI has long been a predominantly technical field, where AI biases were primarily addressed through technical solutions (e.g., Deshpande2020Mitigating; Ramaswamy2021Fair; Ranaldi2023Trip; Zhou2022Bias). As AI bias was identified as a critical issue, the field began examining its effects on human decision-making (e.g., de-arteaga_case_2020; yuan_contextualizing_2023). More recent work has started to incorporate cognitive biases into this context (Kliegr2021review; Bertrand2022How; liao2021xai). Much of the work to date comprises literature reviews that connect insights from other disciplines to AI biases and propose possible solutions (Boonprakong2025cogbiases; Kliegr2021review; Bertrand2022How; Wang2019Designing). Most studies use dual-process theory to broadly categorize cognitive biases as system one thinking (fast and intuitive) or system two thinking (slow and deliberate) (bucinca_trust_2021; lu_human_2021; Wang2019Designing). However, few have directly investigated how specific human biases influence AI-assisted decision-making, lacking specificity in naming cognitive biases in HCI was also criticized in a recent review by Boonprakong2025cogbiases.
@@
Cognitive biases, including BRN, have been linked to AI interpretation, with proposed de-biasing strategies (Kliegr2021review). Although empirical research on the role of cognitive biases in human-AI interaction remains sparse, recent studies are beginning to address this gap. The influence of cognitive biases in human-AI interaction has been empirically examined, with biases broadly operationalized as ’system one’ thinking (lu_human_2021; bucinca_trust_2021)). More specific investigations demonstrated the impact of anchoring bias, showing that positive first impressions of AI can lead to increased over-reliance (nourani_anchoring_2021). Following this, some algorithmic strategies for mitigating anchoring bias were proposed (Echterhoff2022anchoring). Research on the Dunning-Kruger effect, a meta-cognitive bias, revealed that individuals who overestimate their abilities tend to under-rely on AI when they are incorrect (He2023Dunning). A recently published study was the first to investigate the amplifying effects of AI and human biases, highlighting feedback loops that can occur across different contexts (glickmanHowHumanAI2024). Although their study was unpublished and thus unavailable to us during the course of our work, our research similarly examines how AI bias interacts with human bias. However, we extend their findings by exploring different types of bias in a different domain. We propose that interactions between AI and human biases can lead to suboptimal outcomes in AI-assisted decision-making. These biases may compound each other, creating feedback loops that reinforce them. Like vonfeltenIsolationInteractionistPerspective2025, we use the term compound human-AI bias instead of ”bias amplification”, as we believe biases may not only amplify each other but also interact in other ways.
@@ 3. Study Overview and Objectives
Our pre-registered within-subject study on Prolific used a prototypical disease screening tool with two models: one trained on balanced data and one on unbalanced data, where dengue cases were oversampled. This defined AI bias due to class imbalance as the predictor variable model condition (balanced vs. unbalanced). An operationalisation which was also used by  datacentric2021Anik.
@@
Following Schemmer2023Appropriate, participants provided an initial estimate before re-evaluating their answers after receiving AI advice, defining the predictor variable advice condition with two levels: pre-advice vs. post-advice. Using participants’ decisions and the ground truth, we then calculated RAIR and RSR as dependent variables, together measuring the appropriateness of reliance (see Measures for details).
After pre-registration, we observed that repeated AI interactions could influence pre-advice estimates, yielding insights beyond mere AI reliance. To capture these temporal dynamics, we added a trial variable, following Glickman and Sharot (glickmanHowHumanAI2024). This allowed us to distinguish between simple AI over-reliance (a possible reason for post-advice changes) and the gradual adoption of AI bias across repeated trials, even before receiving advice.
We used this screening tool to investigate how AI and human biases interact. Such tools are increasingly relevant as healthcare systems face rising demand and specialist shortages. The use of lay participants reflects real-world contexts: self-diagnosis tools are actively considered for dengue (husinEarlySelfdiagnosisDengue2018), and distinguishing between the common cold and influenza remains a well-documented challenge (mayrhuberFeverItsReal2018). Differentiating among all three diseases is also a problem AI is increasingly applied to solve (nadda_influenza_2022; husinEarlySelfdiagnosisDengue2018). Since early symptoms of dengue closely resemble those of influenza and other viral illnesses, accurate differentiation by lay users is particularly critical (Chambers2019Digital; malavigeDifferentiatingDengueOther2023).

### v2

@@ Abstract.
Figure 1. Conceptual illustration of AI class imbalance, its implications for appropriate reliance calibration and compounding effects with human base rate neglect.
A schematic figure with two panels. Left panel shows the effects of AI Bias on Reliance Calibration. Class imbalance (balanced vs. unbalanced) is shown feeding into an AI systems. A human head in the centre connects AI output to reliance outcomes, signifying human interaction with AI. The balanced input leads to appropriate human reliance, represented by a target with tightly clustered hits. Unbalanced input leads to inappropriate reliance, represented by a scattered target pattern. Right panel shows Compound Human-AI Bias. Human bias (base rate neglect) combines with AI bias (class imbalance). Three pathways are shown: Human with bias alone, unbalanced AI and human bias, and balanced AI and human bias. Each pathway connects to a human head and then to scales comparing human judgment and AI influence. The first two pathways show distorted weighting, with the human-AI combination being the worst, while the balanced pathway shows an equilibrium.
@@ 1. Introduction
Artificial intelligence (AI) is increasingly present in daily life, raising concerns about biased outputs and their societal implications (O’neil2016Weapons; Mallari et al., 2020). In the human-computer interaction (HCI) community, there is a large body of research trying to make AI less biased and therefore more reliable, fair, and transparent (Adadi and Berrada, 2018; Lepri et al., 2018). At the same time, we have a long and deep understanding that human decision-making is susceptible to cognitive biases, causing decisions to deviate from rationality (Tversky and Kahneman, 1974; Wilke and Mata, 2012). As human-AI interaction becomes more common, it is critical to understand what happens when biased AI decision-support meets biased human decision-making. This interplay between human and AI biases has been termed compound human-AI bias (von Felten, 2025).
@@
One possible outcome of such compounding is bias amplification, as observed in areas like emotion perception, motion discrimination, and social judgments (Glickman and Sharot, 2024), or investment decisions (Winder et al., 2025). However, while bias amplification has been studied, the broader dynamics of bias interactions are still not well understood. In particular, the absence of AI bias may also mitigate human bias, suggesting that AI systems have the potential not only to exacerbate existing prejudices but also to influence human decision-making in a positive direction. Understanding these interactions requires examining how human and AI biases influence one another, as well as the conditions under which AI may serve as a corrective mechanism rather than a reinforcing agent of bias. In human-AI interaction, decisions often manifest in reliance, making it a critical outcome for examining how human and AI biases jointly shape behaviour.
@@
Reliance can be defined as a “user’s behavior that follows from the advice of the system” (Scharowski et al., 2022, p. 3). It must be calibrated to AI’s performance to ensure appropriate reliance. Appropriate reliance means relying on AI when it is correct and disregarding it when it is incorrect  (Lee and See, 2004; Schemmer et al., 2023). However, AI bias could complicate and distort this calibration. A common source of bias in AI is class imbalance, which occurs when certain outcome variables are under-represented in AI training data (Fazelpour and Danks, 2021). This imbalance leads to AI systems favouring the majority class, reducing prediction accuracy on new data. While AI bias can skew the AI’s predictions, human biases could further complicate AI-assisted decision-making.
@@
A common human bias, base rate neglect (BRN), can be defined as the tendency to underweight base rate information when estimating probabilities (Yang and Wu, 2020). It poses challenges in decision-making, particularly in AI-assisted tools (Babic et al., 2021). Although many AI tools are developed specifically for experts (Panigutti et al., 2022), non-expert users frequently rely on applications such as symptom checkers or disease screening tools (Cheng et al., 2019), where BRN is important (Babic et al., 2021). Erroneous decisions caused by biases, such as BRN, could increase user distress (White and Horvitz, 2009) and translate to systemic outcomes like increased healthcare costs (Chambers et al., 2019). Despite research on technical AI bias mitigation (Deshpande et al., 2020; Ramaswamy et al., 2021; Ranaldi et al., 2024; Zhou et al., 2022), as well as a plethora of research on cognitive bias mitigation (Korteling et al., 2021), little attention has been given to the biases of humans when interacting with AI. Human-centred AI research is only recently highlighting the importance of cognitive biases in human-AI interactions (Boonprakong et al., 2025; Kliegr et al., 2021; Bertrand et al., 2022; Wang et al., 2019; Liao and Varshney, 2021). However, experimental studies on these interactions are scarce, and the compound effects of AI and human biases remain underexplored.
@@ 2.1. AI Biases: Class Imbalance
Algorithmic bias, or AI bias, refers to systematic deviations in AI outputs caused by biased data at different stages of development, such as data collection, modelling, or deployment (Fazelpour and Danks, 2021). Since AI learns from patterns in its training data, biased data can lead to biased outcomes. These biases may stem from societal inequalities (e.g., gender bias in medical diagnoses (Al Hamid et al., 2024)) or from decisions during data collection that create unrepresentative datasets (Fazelpour and Danks, 2021). For instance, training a model on data that underdiagnoses heart disease in women could lead to male-centric predictions, ultimately reinforcing healthcare disparities (Al Hamid et al., 2024). A common form of data bias is sampling bias, where certain groups are underrepresented, leading to predictions that reinforce existing imbalances (Srinivasan and Chander, 2021). Consider a disease screening tool trained on data collected during a dengue outbreak, where dengue cases were more prevalent than influenza or the common cold. Due to this class imbalance, the tool may later overdiagnose dengue while neglecting the more common diseases.
@@ 2.2. Human-AI Decision-Making: Reliance
Figure 2. Illustration of reliance outcome Combinatorics. This figure was adapted from Schemmer et al. (2023).
This flowchart illustrates how human decisions interact with AI advice. It shows a sequence of initial human decisions, the AI’s correct or incorrect advice, and the final human decision after considering the AI advice. Resulting in eight different reliance outcomes, of which two are correct reliance, and two are incorrect reliance.
Human-AI decision-making, also referred to as AI-assisted decision-making, refers to a collaborative process between humans and AI where AI advice supports users in forming final judgments or decisions (Chen et al., 2023). A common objective in this context is determining whether the AI advice helps users make better decisions (e.g., Nourani et al. (2021); Liu et al. (2021); Zhang et al. (2022)). For instance, an AI-based screening tool for diagnosing dengue fever cases should ideally improve diagnostic accuracy beyond the user’s judgment or the AI’s advice alone.
In this collaborative process, reliance is a critical concept for assessing the effectiveness of human-AI interaction. Reliance can be quantified by assessing how often or how much users adjust their behaviour after receiving AI advice. However, reliance is only beneficial if it is calibrated to the AI’s capabilities (Lee and See, 2004). Ideally, users should demonstrate appropriate reliance (Schemmer et al., 2023; He et al., 2023), relying on the AI when it is correct and refraining from relying when it is incorrect. To measure reliance in these scenarios, a step-wise evaluation of AI advice in the presence of systematic errors is necessary. Collaborative decisions must be compared against uninfluenced human decisions as a baseline. Following these considerations, Schemmer et al. (2023) proposed a measurement concept for classification problems to assess the appropriateness of reliance, distinguishing four reliance outcomes:
@@
Appropriateness of Reliance is captured by two measures: relative self-reliance (RSR) and relative AI-reliance (RAIR). Optimally, both would equal 1, meaning users consistently rely on themselves when they are correct, and the AI is not (RSR=1), and on the AI when it is correct, and they are not (RAIR=1). Together, these measures form a two-dimensional framework, which enables us to test how AI bias affects dimensions of appropriate reliance. However, beyond appropriate reliance, AI bias may also temporally interact with human biases when shaping decisions. If users interact with biased AI, their own biases may compound errors made by the AI, creating a feedback loop (Glickman and Sharot, 2024).
@@ 2.3. Human Biases: Base Rate Neglect
To understand the phenomenon of BRN, a fundamental understanding of Bayes’ theorem is essential. Bayes’ theorem helps to understand how probabilities change after observing new evidence. For example, it helps us determine the probability that a patient has a disease after a positive test result. More formally, it calculates the conditional probability of a disease $H_{\text{disease}}$ given the data $D_{\text{test positive}}$ (for a comprehensive explanation see (Webb and Sidebotham, 2020)). A well-known example illustrating the importance of base rates is regular breast cancer screening for women over 40 (Eddy, 1982). Even with mammograms showing 80% sensitivity and 90% specificity, the low base rate of breast cancer (1%) means that only about 7% of women who test positive once actually have cancer. This probability, known as the positive predictive value (PPV), often surprises people, including physicians, who mistakenly report sensitivity or the difference between sensitivity and specificity as PPV (Gigerenzer et al., 1998; Hoffrage and Gigerenzer, 1998). Because of this neglect of base rates, the usefulness of such screenings is still debated in contemporary medical literature (Jørgensen2016Breast). Traditionally, experiments have focused on assessing BRN through deviations from PPV, requiring participants to integrate base rates with sensitivity and specificity (Dunwoody et al., 2005). However, this approach may even evoke BRN by overwhelming participants with complex statistical concepts (Todd and Goodie, 2002). Patients were shown to mistake sensitivity for PPV in common diseases like strep throat and heart attacks (Hamm and Smith, 1998), and even expert clinicians face challenges with these judgments (Gigerenzer et al., 1998, 2007). To address challenges in probability judgments, the use of natural frequencies (e.g., one out of 1000) has been proposed as an alternative to relative frequencies (e.g., 0.1%) to reduce complexity and mitigate BRN (Hoffrage, 2002). However, even natural frequency formats still require some mental calculations. We argue that to assess BRN with less complexity for participants, presenting them verbally (e.g., ”very often,” ”rarely”) could be more effective. This eliminates the need for mental calculations by participants but still allows to infer BRN by comparing decisions between low- and high-base rate disease. A further advantage is that it does not require participants to know the true population base rate.
@@ 2.4. Human-AI Biases in AI-Assisted Decision-Making
AI has long been a predominantly technical field, where AI biases were primarily addressed through technical solutions (e.g., Deshpande et al. (2020); Ramaswamy et al. (2021); Ranaldi et al. (2024); Zhou et al. (2022)). As AI bias was identified as a critical issue, the field began examining its effects on human decision-making (e.g., De-Arteaga et al. (2020); Yuan et al. (2023)). More recent work has started to incorporate cognitive biases into this context (Kliegr et al., 2021; Bertrand et al., 2022; Liao and Varshney, 2021). Much of the work to date comprises literature reviews that connect insights from other disciplines to AI biases and propose possible solutions (Boonprakong et al., 2025; Kliegr et al., 2021; Bertrand et al., 2022; Wang et al., 2019). Most studies use dual-process theory to broadly categorize cognitive biases as system one thinking (fast and intuitive) or system two thinking (slow and deliberate) (Buçinca et al., 2021; Lu and Yin, 2021; Wang et al., 2019). However, few have directly investigated how specific human biases influence AI-assisted decision-making, lacking specificity in naming cognitive biases in HCI was also criticized in a recent review by Boonprakong et al. (2025).
@@
Cognitive biases, including BRN, have been linked to AI interpretation, with proposed de-biasing strategies (Kliegr et al., 2021). Although empirical research on the role of cognitive biases in human-AI interaction remains sparse, recent studies are beginning to address this gap. The influence of cognitive biases in human-AI interaction has been empirically examined, with biases broadly operationalized as ’system one’ thinking (Lu and Yin, 2021; Buçinca et al., 2021)). More specific investigations demonstrated the impact of anchoring bias, showing that positive first impressions of AI can lead to increased over-reliance (Nourani et al., 2021). Following this, some algorithmic strategies for mitigating anchoring bias were proposed (Echterhoff et al., 2022). Research on the Dunning-Kruger effect, a meta-cognitive bias, revealed that individuals who overestimate their abilities tend to under-rely on AI when they are incorrect (He et al., 2023). A recently published study was the first to investigate the amplifying effects of AI and human biases, highlighting feedback loops that can occur across different contexts (Glickman and Sharot, 2024). Although their study was unpublished and thus unavailable to us during the course of our work, our research similarly examines how AI bias interacts with human bias. However, we extend their findings by exploring different types of bias in a different domain. We propose that interactions between AI and human biases can lead to suboptimal outcomes in AI-assisted decision-making. These biases may compound each other, creating feedback loops that reinforce them. Like von Felten (2025), we use the term compound human-AI bias instead of ”bias amplification”, as we believe biases may not only amplify each other but also interact in other ways.
@@ 3. Study Overview and Objectives
Our pre-registered within-subject study on Prolific used a prototypical disease screening tool with two models: one trained on balanced data and one on unbalanced data, where dengue cases were oversampled. This defined AI bias due to class imbalance as the predictor variable model condition (balanced vs. unbalanced). An operationalisation which was also used by  Anik and Bunt (2021).
@@
Following Schemmer et al. (2023), participants provided an initial estimate before re-evaluating their answers after receiving AI advice, defining the predictor variable advice condition with two levels: pre-advice vs. post-advice. Using participants’ decisions and the ground truth, we then calculated RAIR and RSR as dependent variables, together measuring the appropriateness of reliance (see Measures for details).
After pre-registration, we observed that repeated AI interactions could influence pre-advice estimates, yielding insights beyond mere AI reliance. To capture these temporal dynamics, we added a trial variable, following Glickman and Sharot (Glickman and Sharot, 2024). This allowed us to distinguish between simple AI over-reliance (a possible reason for post-advice changes) and the gradual adoption of AI bias across repeated trials, even before receiving advice.
We used this screening tool to investigate how AI and human biases interact. Such tools are increasingly relevant as healthcare systems face rising demand and specialist shortages. The use of lay participants reflects real-world contexts: self-diagnosis tools are actively considered for dengue (Husin et al., 2018), and distinguishing between the common cold and influenza remains a well-documented challenge (Mayrhuber et al., 2018). Differentiating among all three diseases is also a problem AI is increasingly applied to solve (Nadda et al., 2022; Husin et al., 2018). Since early symptoms of dengue closely resemble those of influenza and other viral illnesses, accurate differentiation by lay users is particularly critical (Chambers et al., 2019; Malavige et al., 2023).